OpenStudy (anonymous):

expand (x-3a)^-3 See my answer in comments

OpenStudy (anonymous):

\[x ^{-3}-3.x ^{-4}.(-3a)+\frac{ -3(-3-1) }{ 2! }x ^{-5}(-3a)^{2}+ \frac{ -3(-3-1)(-3-2) }{ 3! }x ^{-6}(-3a)^{3}\]

OpenStudy (anonymous):

\[x ^{-3}+9x ^{-4}a+54x ^{-5}a ^{2}+270x ^{-6}a\]

OpenStudy (anonymous):

I get all positive answers, yet i know that every second term ie \[9x ^{-4}a\] and \[270x ^{-6}a ^{3}\] are supposed to be negative.

OpenStudy (anonymous):

any suggestions as to why I am an idiot?

hartnn (hartnn):

i would have used the formula... \((a-b)^3 = a^3-3a^2b+3ab^2-b^3\) unless you have typed your question incorrectly...

hartnn (hartnn):

a=x,b=3a

OpenStudy (anonymous):

no its supposed to be to power -3. I have to use binomial theorem

OpenStudy (anonymous):

n = -3

OpenStudy (anonymous):

blast. i typed it wrong in first post

hartnn (hartnn):

so, why did you take 2nd term = (-3a)

hartnn (hartnn):

2nd term wil be 3a if you are using formula which contain - in middle

hartnn (hartnn):

2nd term wil be -3a if you are using formula which contain + in middle

OpenStudy (anonymous):

the binomial theorem states \[(x+b)^{v} = x ^{n}+nx^{n-1}b+\frac{ n(n-1) }{ 2! }x^{n-2}b^{2}+\frac{ n(n-1)(n-2) }{ 3! }x^{n-3}b^{3}+...+b^{n}\]

hartnn (hartnn):

yes, so your 2nd term = -3a and all signs in between will be + yet you've taken '-' in between

hartnn (hartnn):

2nd term = +3x^{-4}(-3a) okk...so you got it now?

hartnn (hartnn):

which you wanted..

OpenStudy (anonymous):

\[-3.x ^{-5}.(-3a) = 9x ^{-4}a\] am I right?

hartnn (hartnn):

why '-' in beginning ??

OpenStudy (anonymous):

as n = -3

OpenStudy (anonymous):

\[(x-3a)^{-3}\] a=x b=-3a n=-3

OpenStudy (anonymous):

\[na ^{n-1}b\]

hartnn (hartnn):

oh wait! all terms positive is correct! who said you are incorrect? :P

OpenStudy (anonymous):

my text book.

hartnn (hartnn):

then there is a typo in your textbook http://www.wolframalpha.com/input/?i=%28x%E2%88%923a%29%5E%28%E2%88%923%29

OpenStudy (anonymous):

Thanks man.

hartnn (hartnn):

welcome ^_^